OpenStudy (anonymous):

solve the equation for 0 less than and equal to x less than 2pi. 4cos^2x-3=0 Thanks.

OpenStudy (anonymous):

cos^2x = 3/4 (cosx)^2 = 3/4 cosx = sqrt(3)/2 Do you know what to do next? Solve for x.

OpenStudy (anonymous):

The equation is : \[4\cos^2x-3=0\] So : \[4\cos^2x=3\] then : \[\cos^2x=\frac34\] so : \[\cos x=\pm\frac{\sqrt3}2\] so the possible values of x are : \[\frac\pi6,~\pi-\frac\pi6,\pi+\frac\pi6, 2\pi-\frac\pi6\]

OpenStudy (anonymous):

thank you :)